OpenStudy (anonymous):

Establish the identity cos^4x = 1/8(3+4cos(2x) + cos (4x))

Parth (parthkohli):

There's really nothing, y'know. Just write \(\cos(4x) = \cos(2 \times 2x)\) and \(\cos(2x) = 2\cos^2 x - 1\).

OpenStudy (anonymous):

i have the answer but i need to know how to do it...

Parth (parthkohli):

They have used standard trig identities, the most important one being \(\cos^2 x = \dfrac{1 + \cos(2x)}{2}\)